For the frame shown, determine the reactions at the pins, namely, A, B, C. 3 m B 1.5 kN/m 1.5 m D 1.5 m 2 kN/m А
Answer #1

3 m BX Consider FBD O AB EMAS O Bx (3) + By (3)-3(1)= 0 BStBy=1 & By *213)=3KN Rm7 AX AyBC By Вх i 0.5)(3)=4.5 KIV 1-5m cx FBD of EMc=0 - Bx (3)+45(15)=0 Bu-2.25 KN EFx=0 ỞxtCr = 5 2:25+ Ga=45 Co=225 KN +1 EFy= 0